OpenStudy (anonymous):

Which of the following is equal to square root of the cube root of 6?

OpenStudy (johnweldon1993):

\[\huge \sqrt{\sqrt[3]{6}} = \sqrt[2\times 3]{6} = \sqrt[6]{6}\]

OpenStudy (anonymous):

A. 6 1/6 B. 6 1/3 C. 6 2/3 D. 6 3/2

OpenStudy (johnweldon1993):

\[\huge \sqrt[6]{6} = 6^{\frac{1}{6}}\]

OpenStudy (anonymous):

thanks a lot!

OpenStudy (johnweldon1993):

No problem!